Projection Admin 3/12
To do this evening/ was completed: Staff screenings have been emailed out, scheduled on the TMS and blocked on the calendar. Had our cinema meeting this morning and we've got everything we need for this upcoming week, sans the slide for They Cloned Tyrone, but other than that, content is all set.
Got my eyes and ears on the shorts Wait to Tell My Mother, that we were having difficulties with. It looks good. We'll take a further look with everyone on Friday when we tech NoBudge. As for the other short, Wire and Cloth, I reached out to the filmmaker and we'll be receiving a prores from them tonight, which can DCP for Friday.
There's a closed captioning device currently in use in T1. Seat 1718.
There's a closed captioning device currently in use in T1. Seat 1718.
Got in a little earlier than usual for the private event this morning, Superhuman Body screening and Q&A, which went very well. FYI - I've made a note about this in the Weekly Update, but please keep this DCP on the server as we'll be showing it again this upcoming week as well as the first week of April. We have the CRU drive in house, but it's just easier to keep it ingested (unless we desperately need storage in T1).
Presentation Issues:
Preshow Clips/Playlist: Continued building out SE playlists for this week.
Trailer Updates:
Special Events/ AV set up: Setup the microphone in T4 (tested it), ran loops and a small reel of Pariah for the show tonight, just to double check focus. It really looks incredible. Should be a great show. Put a new battery in the T4 spots remote. All set.
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): Ingested Snow White this morning.
Presentation Issues:
Preshow Clips/Playlist: Continued building out SE playlists for this week.
Trailer Updates:
Special Events/ AV set up: Setup the microphone in T4 (tested it), ran loops and a small reel of Pariah for the show tonight, just to double check focus. It really looks incredible. Should be a great show. Put a new battery in the T4 spots remote. All set.
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): Ingested Snow White this morning.
Paper Moon came in via satellite this afternoon, but it's not available on the Deulxe server just yet. I'd check back in a little later this evening.
Outgoing/Shipping (content):
Print/video Inspection:
Asset tracking doc updated/checked?:
Maintenance/cleaning: T3 briefly went offline this morning. I reseated the board and that did the trick, but this, along with a few other maintenance bits, have been passed on to DMS.
Cue sheets updated/checked (every shift): Yes.
Marquee updated?:
Posters updated?:
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Comments
Post a Comment